• Log InLog In
  • Register
Liquid`
Team Liquid Liquipedia
EST 18:41
CET 00:41
KST 08:41
  • Home
  • Forum
  • Calendar
  • Streams
  • Liquipedia
  • Features
  • Store
  • EPT
  • TL+
  • StarCraft 2
  • Brood War
  • Smash
  • Heroes
  • Counter-Strike
  • Overwatch
  • Liquibet
  • Fantasy StarCraft
  • TLPD
  • StarCraft 2
  • Brood War
  • Blogs
Forum Sidebar
Events/Features
News
Featured News
[ASL22] Finals Preview: Sweet Victory!12[ASL22] Ro4 Preview: Mirror Mirror6[ASL22] Ro8 Preview: Within Reach5[ASL22] Ro8 Preview: In A Tizzy11[ASL22] Ro16 Preview: Holy Diver5
Community News
Weekly Cups (Oct 26 - Nov 1): MaxPax, herO, Solar win0Weekly Cups (October 19-25): Clem triples first in 5.0.175Evil Geniuses - shutting down23ISSL (IntoTheTV x SOOP SC2 League): Premier7Weekly Cups (October 12-18): 5.0.17 era begins1
StarCraft 2
General
[D] Hacker peak in sc2 ? No more SC2 streamers on team liquid? Liquibets sc2 season 30 Evil Geniuses - shutting down Weekly Cups (Oct 26 - Nov 1): MaxPax, herO, Solar win
Tourneys
SC4ALL: II Venue Change Announcement! Sparkling Tuna Cup - Weekly Open Tournament WardiTV Mondays ISSL (IntoTheTV x SOOP SC2 League): Premier Stellar Fest TWO the Moon (Dec 16-20)
Strategy
[H] ZvP Mid-Late Game: Stalkers Collossi HT
Custom Maps
[G] How to playtest arcade maps Starcraft 2 Advance [D] the control group Simple Questions/Answers
External Content
Mutation # 550 Graveyard Shift The PondCast: SC2 News & Results Mutation # 549 Endless Sparkles Mutation # 548 Shining Bright
Brood War
General
BGH Auto Balance -> http://bghmmr.eu/ Rollback Netcode in Brood War SC4ALL: II Venue Change Announcement! Ziemie's Starcraft 1 Prequel Campaign 🇺🇦 Ukrainian StarCraft:BroodWar community reborn
Tourneys
[$1250] International ShieldBattery League [Megathread] Daily Proleagues Small VOD Thread 2.0 KCM Race Survival 2026 Season 4
Strategy
50,000-Word Terran Builds Memo 2026-10-22 Cliff Jump Revisited (1 in a 1000 strategy) Simple Questions, Simple Answers
Other Games
General Games
Nintendo Switch Thread Stormgate/Frost Giant Megathread Deus Ex 3!? Mobile Legends: Bang Bang Armies of Exigo - YesYes?
Dota 2
Dota 2 Champions League Season 3 Begins April 25!
League of Legends
[TL LoL EUW IHs] Teemo shall perish
Heroes of the Storm
Heroes of the Storm 2.0
Hearthstone
Deck construction bug
TL Mafia
TL Mafia Community Thread
Community
General
US Politics Mega-thread Dating: How's your luck? Trading/Investing Thread Things Aren’t Peaceful in Palestine Canadian Politics Mega-thread
Fan Clubs
Fan Club Resource Thread Reynor Fan Club The Flash Fan Club
Media & Entertainment
Anime Discussion Thread Movie Discussion! [Manga] One Piece
Sports
Football (Soccer) Thread TeamLiquid Health and Fitness Initiative For 2023 MLB/Baseball 2023
World Cup 2022
Tech Support
Computer Build, Upgrade & Buying Resource Thread
TL Community
Recent Gifted Posts
Blogs
Why I built Repalyzer (1.1.…
BlueStar
Innovate or Game Over?
TrAiDoS
ShieldBattery on Mac
Kraekkling
[ASL22] Ro4 Day1 Ticket Giv…
bITt.mAN
Can Bots Beat Pros?? Starcr…
namkraft
Customize Sidebar...

Website Feedback

Closed Threads



Active: 8263 users

D2L betting: Length of series

Blogs > motbob
Post a Reply
motbob
Profile Blog Joined July 2008
United States12546 Posts
Last Edited: 2014-09-08 03:32:03
September 08 2014 00:41 GMT
#1
If you bet regularly on Dota2lounge, you are probably familiar with the fact that you need to take the length of a series into account. As the series gets longer, you should be willing to take odds that are more skewed towards the favorite. An example: today there was a match between Empire and Alliance. Empire has looked significantly better than Alliance recently, so the 60/40 odds look off at first glance... until you realize the match is a Bo1. OK, that's fine, Bo1 games are pretty high variance. The odds would probably be something like 65/35 if it were a Bo3.

But I worked out something interesting regarding the so-called "best of two" format where teams teams play two games (with the potential of a 1-1 tie). It turns out that, on Dota2lounge, you should treat Bo2 matches the same as a Bo5 for betting purposes. The favored team is about as likely to win you rares in a Bo2 as it would be to win a Bo5. Here's the math comparing a Bo3 and a Bo2:

Let's say Team A has a 60% chance of beating Team B in a given game.

In a Bo2, Team A has a .6*.6=.36 chance of going 2-0, while Team B has a .4*.4=.16 chance. The chance of a 1-1 is 0.6*0.4*2=0.48, but Dota2lounge returns your items in the case of a tie, so we need to focus only on the two cases where a team goes 2-0. Items are won/lost 52% of the time (.36+.16), and Team A goes 2-0 36% of the time, so we just take 36%/52%=0.69 to find out the proper odds for betting on Team A (69/31).

In a Bo3, Team A has a .36 chance of going 2-0. As mentioned above, the chance of a 1-1 is 0.48, and Team A has a 60% chance of winning the third game, so the chance of Team A winning in three games is .48*.6=0.288. Since either going 2-0 or winning the third game are the only two ways Team A can win the series, we add .36 and .288 to get the chance of Team A winning the series (0.65, so the proper odds for a Bo3 are 65/35).

69% for a Bo2 vs 65% for a Bo3. I won't do the math here, but the chance of Team A winning a Bo5 against Team B is also about 69%.

If you take issue with my math, I will be happy to defend it to the death in the thread below.
ModeratorGood content always wins.
Dubzex
Profile Joined October 2010
United States6994 Posts
September 08 2014 01:05 GMT
#2
Thoughts on fnatic vs NVMI?
"DONT UNDERESTIMATE MY CARRY OR YOU WILL BE CARRIED INTO THE ABYSS OF SUFFERING" - Tyler 'TC' Cook
motbob
Profile Blog Joined July 2008
United States12546 Posts
Last Edited: 2014-09-08 01:39:23
September 08 2014 01:35 GMT
#3
On September 08 2014 10:05 Dubzex wrote:
Thoughts on fnatic vs NVMI?

Well, the odds are 67/33 right now. According to the OP, I should bet on Fnatic if I think they have about a 59% or greater chance of winning any given game.

I haven't watched Fnatic's games with the new roster, so I dunno. Looking at the two team's results since TI4, I think a bet on Fnatic might be justified but I am not personally going to bet on them having not seen their games.

EDIT: Actually, current key odds are 60/40 which I would judge to be worthy of a bet on Fnatic based on results alone.
ModeratorGood content always wins.
icystorage
Profile Blog Joined November 2008
Jollibee19350 Posts
September 08 2014 02:40 GMT
#4
hi motbob, i made this thread: http://www.liquiddota.com/forum/dota-2-general/454464-dota-2-betting-general

do you mind if i update the OP with your OP?
LiquidDota StaffAre you ready for a Miracle-? We are! The International 2017 Champions!
motbob
Profile Blog Joined July 2008
United States12546 Posts
September 08 2014 03:30 GMT
#5
On September 08 2014 11:40 icystorage wrote:
hi motbob, i made this thread: http://www.liquiddota.com/forum/dota-2-general/454464-dota-2-betting-general

do you mind if i update the OP with your OP?

You can do it if you want. It's up to you if you want to circulate this knowledge.
ModeratorGood content always wins.
icystorage
Profile Blog Joined November 2008
Jollibee19350 Posts
September 08 2014 03:35 GMT
#6
Thanks!
LiquidDota StaffAre you ready for a Miracle-? We are! The International 2017 Champions!
Archeon
Profile Joined May 2011
3274 Posts
Last Edited: 2014-09-08 04:42:22
September 08 2014 04:39 GMT
#7
Interesting, pretty cool. Did the math for some higher winrates to see if it stays true and apparently even 0.99% winrate has a worse chance of loosing at a best of two than a best of three. Which is somewhat counter-intuitive.

Might be worth noting that everything over 2/3 winrate means that the stronger team wins 2:0 is the most probable event.
low gravity, yes-yes!
Rainling
Profile Joined June 2011
United States456 Posts
September 08 2014 20:06 GMT
#8
I can confirm this works, I posted a reddit thread about this a few months ago. It's worth pointing out that the potential losses and gains are both lower in a Bo2 bet because of the possibility of a 1-1 tie, so you should bet a bit more stuff (the more even the matchup is, the more you should bet) on Bo2s relative to Bo5s to get the same expected value.
FiWiFaKi
Profile Blog Joined February 2009
Canada9859 Posts
Last Edited: 2014-09-09 23:16:53
September 09 2014 23:12 GMT
#9
Shoot, I'm trying to think of how to solve this using statistics, and I feel really stupid right now, but I cannot remember.

My question is, what is the chance of the favored team winning a Bo7, if they have a 60% chance to win an individual game.

[image loading]

Using the binomial distribution formula above, I would sub in x = 4 (as that is the victory condition), and take the sum of subbing in n=4,5,6,7, with the p value being 0.6 obviously. I'm getting wonky results though, and negative binomial distribution doesn't seem correct either (probably because it's used for something completely different).

Going through the distributions I've been taught: beta, chi-square, t, f, error, gamma, weibull, poisson, exponential (special case of gamma), or uniform... Definitely none of those are used to solve the problem. Any help would be much appreciated, it's bugging me too much. (I'm not trying to solve this with the brute force method, as that is quite easy [but time consuming] with a little branched tree giving all possible events)

In life, the journey is more satisfying than the destination. || .::Entrepreneurship::. Living a few years of your life like most people won't, so that you can spend the rest of your life like most people can't || Mechanical Engineering & Economics Major
LSB
Profile Blog Joined April 2010
United States5171 Posts
Last Edited: 2014-09-10 18:27:32
September 10 2014 00:06 GMT
#10
1 - (.4^3 + .4^3 * .6 * 3 + .4^3 * .6^2 * 6)

Edit: 68.256%
Once is an accident. Twice is coincidence. Three times is an enemy action. Bus Driver can never target themselves I'm sorry
FiWiFaKi
Profile Blog Joined February 2009
Canada9859 Posts
Last Edited: 2014-09-10 01:44:24
September 10 2014 01:40 GMT
#11
But what did you derive that formula from?

I'm looking for a general solution that I can apply for any situation, not how to solve this problem, that's not my intention.
In life, the journey is more satisfying than the destination. || .::Entrepreneurship::. Living a few years of your life like most people won't, so that you can spend the rest of your life like most people can't || Mechanical Engineering & Economics Major
RuiBarbO
Profile Blog Joined August 2012
United States1340 Posts
September 10 2014 02:10 GMT
#12
Interesting... The math makes sense to me, not that I know much, but I'm wondering: is it really worthwhile to bet on the favored team in a Bo2 series when the odds (before recalculating using motbob's technique) are 60% in their favor? Because while the new odds are quite favorable (69%), that's only looking at scenarios in which only one team wins any games. If the odds of the teams tying is 48%, then the odds of either the teams tying OR the less favored team winning should be 64%, right? So shouldn't this make Bo2s a low-risk environment in which to bet on the less-favored team?
Can someone please explain/how water falls with no rain?
LSB
Profile Blog Joined April 2010
United States5171 Posts
Last Edited: 2014-09-10 18:31:12
September 10 2014 18:29 GMT
#13
On September 10 2014 10:40 FiWiFaKi wrote:
But what did you derive that formula from?

I'm looking for a general solution that I can apply for any situation, not how to solve this problem, that's not my intention.

If you want a general solution just look up Binomial CDF, most statistical software and graphing calculators have the function. Here's the wikipedia link for the equation http://en.wikipedia.org/wiki/Binomial_distribution#Cumulative_distribution_function which was what you were doing. There is no 'one step' equation that replicates the Binomial CDF.

For the specific problem, take 100% - the probability of 0-3, - the probability of 1-3 (times the ways it can happen, 3 ways) - the probability of 2-3 (times the way it can happen)
Once is an accident. Twice is coincidence. Three times is an enemy action. Bus Driver can never target themselves I'm sorry
Rainling
Profile Joined June 2011
United States456 Posts
Last Edited: 2014-09-10 20:13:28
September 10 2014 20:13 GMT
#14
+ Show Spoiler +
On September 10 2014 08:12 FiWiFaKi wrote:
Shoot, I'm trying to think of how to solve this using statistics, and I feel really stupid right now, but I cannot remember.

My question is, what is the chance of the favored team winning a Bo7, if they have a 60% chance to win an individual game.

[image loading]

Using the binomial distribution formula above, I would sub in x = 4 (as that is the victory condition), and take the sum of subbing in n=4,5,6,7, with the p value being 0.6 obviously. I'm getting wonky results though, and negative binomial distribution doesn't seem correct either (probably because it's used for something completely different).

Going through the distributions I've been taught: beta, chi-square, t, f, error, gamma, weibull, poisson, exponential (special case of gamma), or uniform... Definitely none of those are used to solve the problem. Any help would be much appreciated, it's bugging me too much. (I'm not trying to solve this with the brute force method, as that is quite easy [but time consuming] with a little branched tree giving all possible events)


Here's a few reasons why this solution doesn't work: First the n = 7 choose 4 solution includes several impossible scenarios (win-win-win-win-loss-loss-loss, win-win-loss-win-win-loss, etc.) for a Bo7. The n = 6 choose 4 and n = 5 choose 4 similarly include impossible results. Additionally, the probability calculation for an n = 7 choose 4 solution assumes impossible solutions for an actual Bo7. You can't have a loss-loss-loss-loss-loss-loss-loss Bo7, but the n = 7 choose 0 solution is assumed to be an alternate possibility for n = 7 and reduces the calculated probability for n = 7 choose 4. Also you can't sum the n = 7, n = 6, n = 5, n = 4 probabilities, because the n = 4 probability (win - win - win - win) is kind of redundant with a n = 5 possibility (win - win - win - win - loss) and so on.

I have no idea how to do this without brute-forcing though, sorry.
FiWiFaKi
Profile Blog Joined February 2009
Canada9859 Posts
Last Edited: 2014-09-11 01:36:26
September 10 2014 21:36 GMT
#15
Yes, I knew it wasn't right, but it seems like you explained well why that is the case, so thank you.

LSB, I know it needs cumulative binomial distribution, and I know how to solve problems with cumulative binomial distribution, but I simply don't know how to solve approach this one.

edit: Yay, I made solutions:

[image loading]
In life, the journey is more satisfying than the destination. || .::Entrepreneurship::. Living a few years of your life like most people won't, so that you can spend the rest of your life like most people can't || Mechanical Engineering & Economics Major
LSB
Profile Blog Joined April 2010
United States5171 Posts
Last Edited: 2014-09-11 14:00:59
September 11 2014 14:00 GMT
#16
It's not 1-(Binomialcdf (7,4)). It is 1- (Binomialcdf (6,3) * 0.4)

The final game of the series is won be the team that is not favored to win, so you can just imagine it as the probability you will get 3 wins in a series of 6 * the probability of winning the final game.
Once is an accident. Twice is coincidence. Three times is an enemy action. Bus Driver can never target themselves I'm sorry
FiWiFaKi
Profile Blog Joined February 2009
Canada9859 Posts
Last Edited: 2014-09-11 18:18:56
September 11 2014 18:15 GMT
#17
On September 11 2014 23:00 LSB wrote:
It's not 1-(Binomialcdf (7,4)). It is 1- (Binomialcdf (6,3) * 0.4)

The final game of the series is won be the team that is not favored to win, so you can just imagine it as the probability you will get 3 wins in a series of 6 * the probability of winning the final game.


a) If you're doing it from the perspective of the favored team (p>0.5) and trying to solve for the favored team: Binomialcdf(7,3) will give you the probability of the favored team winning 0,1,2, or 3, games, and that is exactly what we don't want, so all we have do to is simply:

P = 1 - binomialcdf(7,3), p=favored

b) If you're doing it from the perspective of the disadvantageous team (p<0.5), but still trying to solve for the favored team: Binomialcdf(7,3), which is the same thing as saying when the disadvantaged team wins 0,1,2, or 3 games, which is the same as the favored team winning, therefore solution is:

P= binomialcdf(7,3), p=disadvantaged

Intuitively this also makes sense, because naturally, we can create this identity, all it's saying is that the sum of probabilities is one:

binomialcdf(7,3)_a +binomialcdf(7,3)_b = 1

So if you have any of the two of the three above equations I sort of derived (but not really), you can get the last one. But in the end, method (a) and method (b) will both net you the same answers.
In life, the journey is more satisfying than the destination. || .::Entrepreneurship::. Living a few years of your life like most people won't, so that you can spend the rest of your life like most people can't || Mechanical Engineering & Economics Major
LSB
Profile Blog Joined April 2010
United States5171 Posts
Last Edited: 2014-09-12 18:25:16
September 12 2014 17:43 GMT
#18
I see what I made a mistake in, for some reason I was calculating best of 5.

For best of 7 it would be
1 - (.4^4 + .4^4 * .6 * 4 + .4^4 * .6^2 * 10 + .4^4 * .6^3 * 20)

But yes your solution is valid
Once is an accident. Twice is coincidence. Three times is an enemy action. Bus Driver can never target themselves I'm sorry
Please log in or register to reply.
Live Events Refresh
Next event in 11h 20m
[ Submit Event ]
Live Streams
Refresh
StarCraft 2
Nathanias 141
StarCraft: Brood War
Britney 7353
Shuttle 375
LuMiX 2
Dota 2
PGG 174
League of Legends
Doublelift4558
JimRising 700
Counter-Strike
jkaem73
Other Games
gofns10329
summit1g8675
Grubby2624
shahzam426
PiGStarcraft371
C9.Mang0268
syndereN165
XaKoH 114
Fuzer 86
NeuroSwarm53
ViBE47
ZombieGrub44
KnowMe26
PPMD22
Railgan21
JuggernautJason4
Maynarde2
Organizations
Other Games
BasetradeTV197
Counter-Strike
PGL83
[ Show 18 non-featured ]
StarCraft 2
• Hupsaiya 48
• RyuSc2 18
• AfreecaTV YouTube
• intothetv
• Kozan
• IndyKCrew
• Laughngamez YouTube
• Migwel
StarCraft: Brood War
• HerbMon 45
• BSLYoutube
• STPLYoutube
• ZZZeroYoutube
Dota 2
• WagamamaTV619
League of Legends
• Nemesis3723
• Jankos3524
• Scarra946
Other Games
• imaqtpie688
• Shiphtur136
Upcoming Events
IntoTheTV X SOOP
11h 20m
Reynor vs Solar
ByuN vs SHIN
Clem vs MaxPax
herO vs Rogue
IPSL
17h 20m
WorsT vs nOmaD
Dewalt vs andersOn
AI Arena Tournament
19h 20m
IntoTheTV X SOOP
1d 11h
IPSL
1d 17h
Azhi_Dahaki vs ggnore13
Double vs WIZARD
Replay Cast
1d 23h
Replay Cast
2 days
Sparkling Tuna Cup
3 days
WardiTV Invitational
3 days
YoungYakov vs Lambo
SHIN vs TBD
Percival vs Shameless
Solar vs TBD
PiGosaur Cup
4 days
[ Show More ]
The PondCast
5 days
KCM Race Survival
5 days
WardiTV Invitational
5 days
Liquipedia Results

Completed

Proleague 2026-11-05
Blizzard Classic Cup 2026
HCC Season 3

Ongoing

Acropolis #5
KCM Race Survival 2026 Season 4
International ShieldBattery League
BSL Season 23
IPSL Winter 2026-27
Calamity Cup Div. A Season 8
IEM Beijing 2026
PGL Masters Bucharest 2026
Stake Ranked Episode 5
1win Private Club #2
Thunderpick World Champ. '26
ESL Pro League Season 24
Stake Ranked Episode 4
1win Private Club #1
Logitech G Play Connect 2026
SL StarSeries Fall 2026
FISSURE Playground #3
BLAST Open Fall 2026

Upcoming

Sector 951
SC4ALL II: Brood War
BSL 23: Non-Korean Championship
HSC XXX
Stellar Fest 2: Lunar Cup
SC4ALL II: StarCraft II
Kung Fu Cup 2026 Grand Finals
RSL Offline Finals
Underdog Cup 2026 Div 2-3
PGL Bucharest 2027
BLAST Bounty Winter 2027
eXTREMESLAND 2026
PGL Major Singapore 2026
Stake Ranked Episode 6
BLAST Rivals Fall 2026
TLPD

1. ByuN
2. TY
3. Dark
4. Solar
5. Stats
6. Nerchio
7. sOs
8. soO
9. INnoVation
10. Elazer
1. Rain
2. Flash
3. EffOrt
4. Last
5. Bisu
6. Soulkey
7. Mini
8. Sharp
Sidebar Settings...

Advertising | Privacy Policy | Terms Of Use | Contact Us

Original banner artwork: Jim Warren
The contents of this webpage are copyright © 2026 TLnet. All Rights Reserved.